OpenStudy (anonymous):

Simplify.\[\Large\tt\frac{8}{5b+2}+\frac{10}{3b-4}\]

OpenStudy (anonymous):

I already know that the denominator would be (5b + 2)(3b - 4). I just need help with the numerator.

OpenStudy (anonymous):

wait i'll solve

OpenStudy (anonymous):

2(37b−6)(3b−4)(5b+2)

OpenStudy (anonymous):

Did you mean: \[\frac{2(37-6)}{(3b-4)(5b+2)}\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

How did you get that answer? :)

OpenStudy (anonymous):

I think I know, but I am not sure. Let me finish what I was doing offline and see if I get the same answer.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can you show me the steps you took to get your answer? @MGLMystiicz

OpenStudy (anonymous):

@rational Can you help me? :)

OpenStudy (rational):

\[\tt\frac{8}{5b+2}+\frac{10}{3b-4}\] You have two fractions with different denominators, so you cannot combine them directly.

OpenStudy (rational):

Get common denominator first

OpenStudy (anonymous):

The common denominator is (5b + 2)(3b - 4).

OpenStudy (rational):

multiply top and bottom of first fraction by 3b-4 multiply top and bottom of second fraction by 5b+2 \[\tt\frac{8(3b-4)}{(5b+2)(3b-4)}+\frac{10(5b+2)}{(3b-4)(5b+2)}\]

OpenStudy (rational):

Now that the bottoms of both fractions are same, you may add the numerators and put the common denominator below them : \[\tt\frac{8(3b-4)+10(5b+2)}{(5b+2)(3b-4)}\]

OpenStudy (rational):

simplify the numerator

OpenStudy (anonymous):

Mhmm. And do we multiply the numerators and then add them? :)

OpenStudy (rational):

yes

OpenStudy (anonymous):

\[\tt\frac{24b-4}{(5b+2)(3b-4)}+\frac{50b+2}{(3b-4)(5b+2)}\] @rational

OpenStudy (rational):

yes, add the numerators and put the common denominator below them

OpenStudy (anonymous):

\[\tt\frac{24b-4+50b+2}{(3b-4)(5b+2)}\]

OpenStudy (anonymous):

@rational ^

OpenStudy (rational):

numerator can be simplified further

OpenStudy (rational):

combine like terms in numerator

OpenStudy (anonymous):

\[\tt\frac{74b-2}{(3b-4)(5b+2)}\]

OpenStudy (anonymous):

@rational ^

OpenStudy (anonymous):

Hey, Mr. @rational, you still there? ._.

OpenStudy (rational):

looks good, also you may factor out "2" from numerator

OpenStudy (anonymous):

How? :)

OpenStudy (rational):

\[\tt\frac{74b-2}{(3b-4)(5b+2)}\] \[\tt\frac{2(37b-1)}{(3b-4)(5b+2)}\]